How much must you deposit each year into your retirement account starting now and continuing through year 11 if you want to be able to withdraw $70,000 per year forever, beginning 28 years from now? Assume the account earns interest at 12% per year.
Answer #1

ANSWER:

I = 12%

We will find the present value of the $70,000 per year 28 years from now on.

PV = Withdrawable amount per year / interest rate (p/f,i,n)

pv = 70,000 / 12% (p/f,12%,27)

pv = 583,333.33 * 0.0469

pv = 27,358.33

Deposit per year = pv(a/p,i,n) = 27,358.33(a/p,12%,11) = 27,358.33 * 0.1684 = 4,607.14

so the deposit per year is $4,607.14